OpenStudy (anonymous):

simplify by removing the factors of 1. 5n^2+5n/10n^2+25n

OpenStudy (callisto):

5n^2+5n/10n^2+25n = 5n(n+1) / 5n (2n+5) = (n+1) / (2n+5)

OpenStudy (anonymous):

thank you